OpenStudy (anonymous):

Can someone help me please?????? What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (luigi0210):

First get rid of y, so multiply the bottom by 2 and the two equations

OpenStudy (anonymous):

is it C?

OpenStudy (luigi0210):

Did you work it out or are you guessing?

OpenStudy (anonymous):

I worked it out but I don't think its right

OpenStudy (luigi0210):

Well don't doubt yourself kiddo, that's right :)

OpenStudy (anonymous):

thank you :)

OpenStudy (luigi0210):

Anytime

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (luigi0210):

Sure

OpenStudy (anonymous):

Rick works two jobs to pay for college. He tutors for $20 per hour and also works as a bag boy for $7 per hour. Due to his class and study schedule, Rick is only able to work up to 25 hours per week, but must earn at least $200 per week. If t represents the number of hours Rick tutors and b represents the number of hours he works as a bag boy, which system of inequalities represents this scenario? t + b less than or greater to 25 20t + 7b greater than or equal to 200 t + b greater than or equal to 25 20t + 7b = 200 t + b less than or greater to 25 20t + 7b less than or greater to 200 None of the systems shown represent this scenario.

OpenStudy (luigi0210):

@whpalmer4 I think it might none but I'll get someone to help

OpenStudy (whpalmer4):

tutoring and bagging time is up to 25 hours per week -> t+b <= 25 the first answer choice, is "less than or greater to" a typo?

OpenStudy (whpalmer4):

same with the 3rd answer choice well, \(t + b \le 25\) is one inequality "he must earn at least $200/week" -> \(20t + 7b \ge 200\)

OpenStudy (anonymous):

less than or equal to

OpenStudy (whpalmer4):

the two inequalities are \[t+b \le 25\]\[20t+7b\ge200\] you decide if that appears on your answer list :-)

OpenStudy (anonymous):

the third choice is < with line underneath

OpenStudy (luigi0210):

Thanks @whpalmer4 :)

OpenStudy (whpalmer4):

you're welcome!

OpenStudy (anonymous):

its A and thank you :)

OpenStudy (whpalmer4):

@Luigi0210 hey, thanks for pushing me to 96 :-)

OpenStudy (luigi0210):

No Problem :D

OpenStudy (whpalmer4):

celebration time, @jaderbrown, do you have another problem? :-)

OpenStudy (anonymous):

yes :) What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

OpenStudy (whpalmer4):

Okay, do you know how to solve by elimination?

OpenStudy (anonymous):

yeah but I always get it wrong or stuck :(

OpenStudy (whpalmer4):

let's see if we can change that :-)

OpenStudy (luigi0210):

@jaderbrown I told you not to doubt yourself >.<

OpenStudy (whpalmer4):

so, which variable shall we eliminate? if we eliminate y, that looks a little easier, but then we need to solve for both x and y. if we eliminate x, then we don't have to solve for a second variable, but the arithmetic is a little uglier. your choice, we can do it both ways, too!

OpenStudy (anonymous):

i know luigi lol, whpalmer can we do it both ways so I can understand it more?

OpenStudy (whpalmer4):

sure. I'll do the ugly one first, then you can do the other one — anyone who saw both of our pictures would agree this is the right division ;-) \[13x - 6y = 20\]\[ 7x + 4y = 18 \]I'm going to eliminate \(x\) by multiplying the first row by 7 and the second row by -13.

OpenStudy (whpalmer4):

as they say at the magic show, watch my hands carefully :-)

OpenStudy (whpalmer4):

\[7*13x - 7*6y = 7*20\]\[-13*7x-13*4y = -13*18\]a bit of arithmetic and then we have \[~~~91x -42y = ~~~140\]\[-91x -52y = -234\]-------------------- add down the columns to get \[(91-91)x +(-42-52)y = 140-234\]\[-94y = -94\]\[y = 1\] Now we use the value of \(y\) in one of the original equations to find the value of \(x\): \[13x-6(1) = 20\]\[13x-6=20\]\[13x=26\]\[x=2\] Check with the other equation: \[7(2)+4(1) = 18\]\[14+4 = 18\checkmark\] [I'm not rechecking with the first equation because I used it to find \(x\) from \(y\)]

OpenStudy (whpalmer4):

The key was just multiplying the first equation by the coefficient of the variable I wanted to eliminate in the second equation, and vice versa. I also multiplied (at the same time) by -1 so that they would be opposite in sign.

OpenStudy (whpalmer4):

For some equations, you already have the coefficients with opposite signs, and then you don't need to do that step.

OpenStudy (whpalmer4):

Okay, now it's your turn. The old surgeon's motto: "see one, do one, teach one" :-)

OpenStudy (anonymous):

didn't you just solve for x and y? so what do I do ?

OpenStudy (whpalmer4):

you're going to do the same thing, except eliminate y instead of x.

OpenStudy (whpalmer4):

you'll get the same answer, of course, breaking the curse that has dogged you :-)

OpenStudy (anonymous):

okay hold on :)

OpenStudy (anonymous):

i got stuck at x=6y+20 divided by 13

OpenStudy (anonymous):

I don't think im doing it right at all. I looked at how you did it and Im not even close to what you did :(

OpenStudy (anonymous):

im suppose to eliminate 13x so i divide it on each side and i divide it from 6y+20 so now what?

OpenStudy (whpalmer4):

let's start at the beginning. \[13x−6y=20\]\[7x+4y=18\] You're planning to eliminate \(y\), and you've already got opposite signs there, right?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

okay, so now we need to do something to those two equations that will get us a y coefficient in the first equation that is -<some number> and the y coefficient in the second equation is +<that same number>

OpenStudy (whpalmer4):

what if you had the following fraction problem to do: \[\frac{1}{6} + \frac{1}{4}\] what would you use for a common denominator?

OpenStudy (whpalmer4):

you'd probably use either 12 or 24, right?

OpenStudy (anonymous):

use 12

OpenStudy (whpalmer4):

okay. what do you need to multiply the first equation by to get the y coefficient to be -12? then, what do you need to multiply the second equation by to get its y coefficient to be +12?

OpenStudy (anonymous):

im confused, what are we doing to suppose to be getting 12 or -12?

OpenStudy (whpalmer4):

we want the y coefficients to cancel out when we add the two equations together - elimination is the name of the game

OpenStudy (anonymous):

okay so you add 6 to 20 so its 26 divide 13 from 26 x=2

OpenStudy (anonymous):

did i do it right?

OpenStudy (whpalmer4):

\[13x−6y=20\]\[7x+4y=18\] what did you multiply with the first equation? what did you multiply with the second equation? (I just want to go through the steps to make sure you've got it all)

OpenStudy (whpalmer4):

\[26x - 12y = 40 \text{ (assuming you multiplied by 2)}\]\[21x+12y=54\text{ (assuming you multiplied by 3)}\]------------------\[47x + 0y = 94\]

OpenStudy (whpalmer4):

the solution to that is clearly \(x =2\) I'm just not quite sure about your "okay so you add 6 to 20 so its 26/ divide 13 from 26" maybe you could show a little more detail?

OpenStudy (anonymous):

well I just did the first equation but do I have to do the second one?

OpenStudy (whpalmer4):

that's entirely up to you. you do already know the answer, but a little practice is always good.

OpenStudy (anonymous):

i think I got it :)

OpenStudy (whpalmer4):

okay, well, I'm off to another problem! good night...